The sum of two consecutive numbers is -61.
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


x = first integer
x+1 = second integer just after x

x+(x+1) = 2x+1 = their sum
2x+1 = -61
2x = -61-1
2x = -62
x = -62/2
x = -31
x+1 = -31+1 = -30

Answer: The two consecutive integers are -31 and -30 in that order.

Answer by ikleyn(52752)

.

Divide -61 by 2 mentally and take two closest integer values to the quotient:

    - one integer number closest from the bottom 

    - and the other integer number closest from the top.